If this reaction produced 51.4 g KCl how many grams of O2 were produced


1
Expert's answer
2021-11-24T03:42:04-0500

Solution:

Calculate the moles of KCl:

The molar mass of KCl is 74.55 g/mol

Therefore,

(51.4 g KCl) × (1 mol KCl / 74.55 g KCl) = 0.6895 mol KCl


Balanced chemical equation:

2KClO3(s) → 2KCl(s) + 3O2(g)

According to stoichiometry:

2 mol of KClO3 produces 2 mol of KCl and 3 mol of O2

Thus, if 0.6895 mol of KCl is produced, then:

(0.6895 mol KCl) × (3 mol O2 / 2 mol KCl) = 1.03425 mol O2

1.03425 mol O2 is produced


Calculate the mass of O2:

The molar mass of O2 is 32.0 g/mol

Therefore,

(1.03425 mol O2) × (32.0 g O2 / 1 mol O2) = 33.096 g O2 = 33.1 g O2


Answer: 33.1 grams of O2 were produced